A lead bullet of mass 50g is fired with velocity 300ms^-1 into a lead block of mass 0.75kg, given that the lead block can move freely, calculate the combine speed of the bullet and the block after the impact


1
Expert's answer
2021-11-14T17:13:56-0500

"mv=(M+m)u,"

"u=v\frac m{M+m}=18.8~\frac ms."